OpenStudy (anonymous):

I need help with this question, The graph below shows the value of Edna's profits f(t), in dollars, after t months: graph of quadratic function f of t having x intercepts at 6, 0 and 18, 0, vertex at 12, negative 36, and passes through point 21, 41.25 What is the closest approximate average rate of change for Edna's profits from the 18th month to the 21st month?

OpenStudy (anonymous):

whees the graph??

OpenStudy (anonymous):

I don't know how to put it on here but the part that says graph of quadratic function f of x that is what is on the graph.

OpenStudy (anonymous):

click on attach file i can't understand u

OpenStudy (anonymous):

Oh ok sorry on my computer the screen is small so i didn't see that okay one sec and i will put the graph.

OpenStudy (anonymous):

ok :3

OpenStudy (anonymous):

Here is the graph

OpenStudy (anonymous):

ok wait up

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

rate of change= slope

OpenStudy (anonymous):

do you know how to find the slope?

OpenStudy (anonymous):

since you are not answering i will show ya

OpenStudy (anonymous):

slope=\[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

so your two points from 18 months to 21 months is: (0,18) and (21,41.25).

OpenStudy (anonymous):

Hey sorry it wouldn't let me type lol

OpenStudy (anonymous):

plug the points into the slope formula and you get:\[\frac{ 41.25-18 }{ 21-0 } =\frac{ 23.25 }{ 21 }\]

OpenStudy (anonymous):

thats the rate o change :3

OpenStudy (anonymous):

g'night btw

OpenStudy (anonymous):

Oh okay thank you! Lol

OpenStudy (anonymous):

:)